Advertisements
Advertisements
प्रश्न
Read the following passage:
|
Recent studies suggest the roughly 12% of the world population is left-handed.
Assuming that P(A) = P(B) = P(C) = P(D) = `1/4` and L denotes the event that child is left-handed. |
Based on the above information, answer the following questions:
- Find `P(L/C)` (1)
- Find `P(overlineL/A)` (1)
- (a) Find `P(A/L)` (2)
OR
(b) Find the probability that a randomly selected child is left-handed given that exactly one of the parents is left-handed. (2)
Advertisements
उत्तर
Given, P(L) = `12/100`
`P(L) P(overlineL) = 1 - 12/100 = 88/100`
and P(A) = P(B) = P(C) = P(D) = `1/4`
`P(L/A) = 24/100`,
`P(L/B) = 22/100`,
`P(L/C) = 17/100`,
`P(L/D) = 9/100`
(i) `P(L/C) = 17/100`, from the given data.
(ii) `P(overlineL/A) = (P(overlineL ∩ A))/(P(A))`
= `(P(A) - P(L ∩ A))/(P(A))`
= `1 - (P(L ∩ A))/(P(A))`
= `1 - P(L/A)`
= `1 - 24/100`
= `(100 - 24)/100`
= `76/100`
= `38/50`
= `19/25`
(iii) (a) `P(A/L) = (P(A ∩ L))/(P(L))`
But `P(L/A) = (P(A ∩ L))/(P(A))`
`24/100 = (P(A ∩ L))/(1/4)`
`\implies` P(A ∩ L) = `24/100 xx 1/4 = 6/100 = 3/50`
∴ `P(A/L) = (3/50)/(12/100) = (3 xx 100)/(12 xx 50) = 1/2`.
OR
(b) `P(L/(B ∪ C)) = (P[(L) ∩ (B ∪ C)])/(P(B ∪ C))`
= `(P[(L ∩ B) ∪ (L ∩ C)])/(P(B ∪ C))`
= `(P(L ∩ B) + P(L ∩ C) - P(L ∩ B)P(L ∩ C))/(P(B) + P(C) - P(B)P(C))` ...(As they are independent)
= `(22/100 xx 1/4 + 17/100 xx 1/4 - 22/400 xx 17/400)/(1/4 + 1/4 - 1/4 xx 1/4)`
= `(22/400 + 17/400 - (22 xx 17)/(400 xx 400))/(1/2 - 1/16)`
= `((39/400 - 374/160000))/(1/2 - 1/16)`
= `16/7((39 xx 400 - 374)/160000)`
= `(16 xx 15226)/(7 xx 160000)`
= 0.217
= 0.22
APPEARS IN
संबंधित प्रश्न
Assume that each born child is equally likely to be a boy or a girl. If a family has two children, what is the conditional probability that both are girls? Given that
- the youngest is a girl.
- at least one is a girl.
The probability that a certain kind of component will survive a check test is 0.6. Find the probability that exactly 2 of the next 4 tested components survive
If P(A) = 0.8, P(B) = 0.5 and P(B|A) = 0.4, find P(A ∪ B)
If `P(A) = 6/11, P(B) = 5/11 "and" P(A ∪ B) = 7/11` find
- P(A ∩ B)
- P(A|B)
- P(B|A)
A fair die is rolled. Consider events E = {1, 3, 5}, F = {2, 3} and G = {2, 3, 4, 5} Find P ((E ∪ F)|G) and P ((E ∩ G)|G)
If P(A) = `1/2`, P(B) = 0, then P(A|B) is ______.
A and B are two events such that P (A) ≠ 0. Find P (B|A), if A is a subset of B.
A die is thrown again and again until three sixes are obtained. Find the probability of obtaining the third six in the sixth throw of the die.
If A and B are events such as that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`, then find
1) P(A / B)
2) P(B / A)
Bag A contains 4 white balls and 3 black balls. While Bag B contains 3 white balls and 5 black balls. Two balls are drawn from Bag A and placed in Bag B. Then, what is the probability of drawing a white ball from Bag B?
Two dice are thrown simultaneously, If at least one of the dice show a number 5, what is the probability that sum of the numbers on two dice is 9?
In an examination, 30% of students have failed in subject I, 20% of the students have failed in subject II and 10% have failed in both subject I and subject II. A student is selected at random, what is the probability that the student has failed in subject I, if it is known that he is failed in subject II?
A bag contains 10 white balls and 15 black balls. Two balls are drawn in succession without replacement. What is the probability that, one is white and other is black?
From a pack of well-shuffled cards, two cards are drawn at random. Find the probability that both the cards are diamonds when the first card drawn is replaced in the pack
Two cards are drawn one after the other from a pack of 52 cards without replacement. What is the probability that both the cards drawn are face cards?
Select the correct option from the given alternatives :
Bag I contains 3 red and 4 black balls while another Bag II contains 5 red and 6 black balls. One ball is drawn at random from one of the bags and it is found to be red. The probability that it was drawn from Bag II
A problem in Mathematics is given to three students whose chances of solving it are `1/3, 1/4` and `1/5`. What is the probability that exactly one of them will solve it?
One bag contains 5 white and 3 black balls. Another bag contains 4 white and 6 black balls. If one ball is drawn from each bag, find the probability that one white and one black
Given P(A) = 0.4 and P(A ∪ B) = 0.7 Find P(B) if A and B are independent events
Choose the correct alternative:
A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2, 5/8`. The probability that the target is hit by A or B but not by C is
A die is thrown nine times. If getting an odd number is considered as a success, then the probability of three successes is ______
If P(A) = `3/10`, P(B) = `2/5` and P(A ∪ B) = `3/5`, then P(B|A) + P(A|B) equals ______.
If two balls are drawn from a bag containing 3 white, 4 black and 5 red balls. Then, the probability that the drawn balls are of different colours is:
A bag contains 3 red and 4 white balls and another bag contains 2 red and 3 white balls. If one ball is drawn from the first bag and 2 balls are drawn from the second bag, then find the probability that all three balls are of the same colour.
Let A and B be two non-null events such that A ⊂ B. Then, which of the following statements is always correct?
Which complement form is correct when \[P(B)\neq 0\]?
For \[E=\{(b,b)\}\] and \[F=\{(b,b),(g,b),(b,g)\}\], what are \[E\cap F\], \[P(E\cap F)\], and \[P(F)\]?
Which statement correctly describes the numerator in \[P(A\mid B)=\frac{P(A\cap B)}{P(B)}\]?

